# Nigel - Mechatronic Design and Robust Sim2Real Control of an Over-Actuated Autonomous Vehicle

### v1

@@ VI Conclusion
In this work, we first introduced the complete mechatronic design architecture of Nigel and also presented the detailed dynamics modeling of this independent 4WD4WS vehicle. We also formulated a linear parameter-varying polytopic model of the system for synthesizing a robust optimal controller using an LMI approach, which seeks to minimize the $H_{2}$-$H_{\infty}$ tradeoff criteria while satisfying the closed-loop pole placement constraints in a Pareto-optimization scheme. We demonstrated and analyzed robust stabilizing as well as steady-state tracking control of Nigel with exhaustive simulation-based verification. The proposed reliable sim2real framework was also experimentally validated using the 4WD4WS vehicle platform. In addition to sim2real transfer, we also validated the controller to handle extra disturbance and uncertainties in real-world conditions.
Future work will delve into formulating and validating a robust tracking problem using mixed sensitivity loop-shaping. Other avenues include fault-tolerant control, and integration of the proposed framework with a high-level autonomy stack, by taking advantage of Nigel’s comprehensive sensor suite.

### v3

@@ VI Conclusion
In this work, we first introduced the complete mechatronic design architecture of Nigel and also presented the detailed dynamics modeling of this independent 4WD4WS vehicle. We also formulated a linear parameter-varying (polytopic) model of the system for synthesizing a robust optimal controller, which seeks to minimize the $H_{2}$-$H_{\infty}$ tradeoff criteria with $D$-stability guarantee. We demonstrated and analyzed robust stabilizing as well as steady-state tracking control of Nigel with extensive experimentation and benchmarking in simulation as well as real world. In addition to sim2real transfer, we also validated the controller to handle exaggerated disturbance and uncertainties in real-world conditions.
Future work will delve into formulating and validating a robust tracking problem using mixed sensitivity loop-shaping. Other avenues include full-scale deployment, fault-tolerant control, and integration of the proposed framework with an end-to-end autonomy stack.